The solubility product constant KS for BaF2 is 1.70*10^–6. How much barium
difluoride can be dissolved in 2.00 liters of pure water?
1
Expert's answer
2012-04-12T09:43:07-0400
s=ks * Mw
s =1.70*10^–6 * 175=2.975*10^-4 g/L
m= 2 L *2.975*10^-4
g/L=0.000595 g.